OpenStudy (anonymous):

how many roots does 2x^2-7x-5=0 have ?

OpenStudy (anonymous):

The discriminant \(\Delta = b^2 - 4ac\) from the quadratic formula \(x = \frac{-b \pm \sqrt{\Delta}}{2a}\)gives us this information. You can see that if delta is negative, there are no real solutions because we cannot take the square root of a negative number; if delta is zero we get one solution (sometimes called a double root) and if it is positive we get two solutions.

OpenStudy (anonymous):

2 should be

OpenStudy (anonymous):

\[ a = 2,\quad b = -7,\quad c = -5 \\ \Delta = (-7)^2 - 4(2)(-5) = 49 + 40 > 0 \] I agree with mohammad.arqum.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome